OpenStudy (pengy120):

When solving the equation for p: q=4p-1. What is the coefficient in front of the q?

OpenStudy (pengy120):

A. -4 B. -¼ C. ¼ D. 4

OpenStudy (whpalmer4):

The coefficient is the number in front. If there is no number, the coefficient is 1.

OpenStudy (whpalmer4):

Now, before you say "but that's not one of my answer choices!" realize that you have to first solve the equation, THEN you look at the coefficient :-)

OpenStudy (pengy120):

Thank you!!!!

OpenStudy (pengy120):

I forgot what coefficent was, thats all..oops

OpenStudy (pengy120):

hello?

OpenStudy (anonymous):

Did you get the answer?

OpenStudy (pengy120):

not yet, they have been typing for a while now..i'm sorta wondering if they are even there

OpenStudy (whpalmer4):

who has been typing for a while now?

OpenStudy (whpalmer4):

can you solve the equation for \(p\)? That means get \(p = \) on the left side, and everything else on the right side...

OpenStudy (anonymous):

Well you have the equation \[q=4p-1\], first you need to solve the equation for p. Do you know how to do that?

OpenStudy (pengy120):

so add the 4p to the left?

OpenStudy (whpalmer4):

well, you can flip the equation over when you're done; it doesn't matter which side you get the variable alone on. Mostly, you want to get everything else off the same side as your variable for which you are solving. \[q = 4p-1\]I might suggest adding 1 to both sides

OpenStudy (anonymous):

\[q = 4p -1\]You need to get p by itself , so add 1 to both sides \[q + 1 = 4p\] Now you need to get p by itself, so divide both sides by 4. What do you get when you divide both sides by 4 ?

OpenStudy (pengy120):

q+1/4=p? im so sorry i have no clue

OpenStudy (anonymous):

That's almost right, except you forgot to divide q by 4, other than that its correct

OpenStudy (anonymous):

\[\frac{ q }{ 4 }+\frac{ 1 }{ 4 } = p\]

OpenStudy (anonymous):

q/4 is that same as \[\frac{ 1 }{ 4 }q\]

OpenStudy (pengy120):

thats what i meant oops

OpenStudy (anonymous):

Oh okay, so you under stand how we got that?

OpenStudy (pengy120):

yes i do

OpenStudy (anonymous):

Great, now to answer the question "What is the coefficient in front of the q?" coefficient is the number in front of the variable. What's in front of q?

OpenStudy (pengy120):

1/4

OpenStudy (anonymous):

That's right!

OpenStudy (pengy120):

so my answer would be C.?

OpenStudy (anonymous):

Yeah

OpenStudy (pengy120):

oh, i had the letter choices above in case you didn't see...well, thank you very much! i think i understamd this now, have a good night/day

OpenStudy (anonymous):

You're welcome